# SubRegWeigh: Effective and Efficient Annotation Weighing with Subword Regularization

### v1

@@ 4.1 Dataset
We briefly overview the dataset used in the experiments here. The detailed information can be found in Appendix B.
@@
We used CoNLL-2003 Tjong Kim Sang and De Meulder (2003). We used the official split for the training, validation, and test datasets, with the label format changed from IOB1 to IOB2. In addition, we used CoNLL-CW Wang et al. (2019) and CoNLL-2020 Liu and Ritter (2023) for the test dataset. [^3] Because both of these dataset names are CoNLL++, we call CoNLL-CW and CoNLL-2020 to distinguish them.
@@
We used SST-2 Socher et al. (2013). Training split was used for the training dataset and validation split was used for the test dataset.
@@ 4.2 Compared Methods
For the baselines, we selected Vanilla and CrossWeigh. Vanilla is the final model that is trained on the original training split without any annotation weighing. CrossWeigh is the final model that is trained with the dataset weighed by the official code [^4] https://github.com/ZihanWangKi/CrossWeigh of the existing work Wang et al. (2019). Although CrossWeigh is specialized to NER, we utilize it for the text classification setting without Entity Disjoint, which excludes sentences containing named entities in the validation data from the training data in each fold. proposed in Wang et al. (2019). This is because Entity Disjoint is unavailable in the text classification.
For the proposed method, we evaluated the three options introduced in §3.2, which are represented as SubRegWeigh (Random), SubRegWeigh (Cos-Sim), and SubRegWeigh ($K$-means), respectively. We used these proposed methods to weight the training dataset and evaluate the final model trained with this weighted dataset as well as other baselines.
@@ 4.3 Model Settings
As the pre-trained language model, we employed $\mathrm{RoBERTa_{LARGE}}$ for both NER and text classification. Besides $\mathrm{LUKE_{LARGE}}$ Yamada et al. (2020) is used for the NER setting because LUKE is an architecture for entity-related tasks. In this experimental setting, we used the same backbone model for both the scouting model and the final model555§5.3 discusses the case where using different pre-trained models between the scouting and final models.. For example, when RoBERTa is used in the scouting model, the final model will also use RoBERTa. The detailed model settings such as hyperparameters are provided in the Appendix C.
CrossWeigh has three hyperparameters: the number of folds in mistake estimation $K$, the number of iterations of mistake estimation $T$, and the weight scaling factor $\epsilon$. We use $K=10$, $T=3$, and $\epsilon=0.7$, respectively. BPE-Dropout Provilkov et al. (2020) was used to obtain the multiple tokenization candidates for SubRegWeigh because both RoBERTa and LUKE employ BPE for their tokenizers. BPE-Dropout has a hyperparameter $p$, where a higher $p$ results in finer segmentation (i.e., more different from the original segmentation). In the experiments, $p$ was set to $0.1$ unless otherwise noted. Additionally, by default, the hyperparameters for SubRegWeigh were set to $N=500$, $K=10$, and $w_{\mathrm{min}}=1/3$.
@@ 4.4 Evaluation Metrics
We evaluate the methods from the viewpoints of speed and performance. On the speed side, we measured the time for detecting the annotation errors and generating the weighted data for the training dataset in NER [^6] We measured the processing time with Xeon Platinum 8167M + NVIDIA V100 for RoBERTa and Xeon Gold 6230R + NVIDIA A100 for LUKE.. On the performance side, we used the weighted data to train the final models and measured the $F_{1}$ score in NER and the accuracy score in text classification for the outputs of the models on each test dataset.
@@
Table 1: The speed of weighing the training dataset and the performance ($F_{1}$) of the final models with the weighted dataset. We measured weighting time with RoBERTa at V100 and LUKE at A100. We report the averaged times and scores and standard deviations over five runs. Vanilla does not have the The best performing scores are highlighted in bold. and †† is SoTA score from Wang et al. (2021b) and  Zhou and Chen (2021).
@@ 4.5 Experimental Results
The results of our experiments are shown in Table 1. We reported the averaged $F_{1}$ scores and accuracy scores over the five independent runs and the standard deviations.
4.5.1 Weighting Time
The column “Weighting time” shows the time taken to detect the annotation errors and generate weights for each sample in the entire training dataset in NER. As shown in this column, the proposed method, SubRegWeigh (Random), was maximum of over eight times faster than CrossWeigh. Even the most time-consuming method of SubRegWeigh ($K$-means) was five times faster with RoBERTa and four times faster with LUKE than CrossWeigh. The method with the random selection is the fastest among the proposed methods because it does not require any of the calculations about TF-IDF, cosine similarity, and $K$-means clustering. In addition, in the random method, only $K(=N=10)$ tokenization candidates were generated, whereas the other methods of SubRegWeigh need to select $K$ candidates from the $N(=500)$-sampled tokenization candidates, which results in longer processing time. Although the proposed method of Cos-Sim and $K$-means takes longer time than Random, it is remarkably faster than the existing method for the automatic annotation weighing, CrossWeigh.
@@ 4.5.2 Performance on NER
For the evaluation of NER, the performance on CoNLL-CW is the most important because annotation errors in this evaluation split are removed, which indicates that we can compare the methods on the most unpolluted dataset. One can see that SubRegWeigh ($K$-means) achieves higher performance than the vanilla method on the test dataset of CoNLL-CW. Specifically, SubRegWeigh ($K$-means) improved the F1 scores by 0.18 points with RoBERTa and 0.2 points with LUKE compared to each vanilla baseline. This result indicates that the proposed method of annotation weighing contributes to the performance improvement in the dataset with fewer annotation errors. Furthermore, SubRegWeigh ($K$-means) achieves higher scores than CrossWeigh, which demonstrates that the proposed method is a reasonable alternative to the existing method.
Both the proposed methods and CrossWeigh scored lower than Vanilla in the CoNLL valid and the original test dataset, which contain annotation errors. These results indicate the remarkable negative effect of the annotation errors in the valid and test datasets and the difficulty to evaluate the models appropriately. The vanilla baseline score of LUKE exceeds the SoTA score Zhou and Chen (2021) for CoNLL-CW, which we consider because of the successful hyperparameter setting.
Among the proposed methods, Random scores lower than $K$-means in many cases. We consider this because the random selection of tokenization candidates leads to the selection of similar subword sequences, causing a bias in the inference results. Cos-Sim also scored lower than $K$-means in all cases, which shows that the naive method of selecting tokenization candidates does not contribute to performance improvement. We consider that one of the reasons for the successful result of the $K$-means selection method is as follows: the $K$-means clustering can handle the diverse range of subword tokenization candidates.
@@ 4.5.3 Performance on Text Classification
For the text classification task, SubRegWeigh ($K$-means) achieved the highest accuracy, and the SubRegWeigh (random) selection had low accuracy, similar to the results in NER. One reason for the lack of performance improvement with CrossWeigh in SST-2 is that Entity Disjoint is important for CrossWeigh but unavailable in the text classification task.
From the experimental results on the NER and text classification datasets, we conclude that the proposed method is superior to the existing method in terms of the weighting time regardless of the methods of selecting tokenization candidates. Especially, with $K$-means, the method achieves approximately four to five times faster performance compared to existing methods, making it highly cost-effective.
@@ 5.1 Pseudo-incorrect Label Test
$\overline{w}_{cor}$ $\overline{w}_{incor}$
@@
Table 2: Averaged weights and performance for the pseudo-incorrect dataset. $\overline{w}_{cor}$ is the averaged weights for the data whose labels were not replaced in the pseudo-incorrect dataset. $\overline{w}_{incor}$ is the averaged weights for the data whose labels were replaced in the pseudo-incorrect dataset.
We verified whether the proposed method can accurately detect annotation errors in the dataset. We evaluate this capability using a modified dataset with some labels flipped to incorrect labels artificially.
Assuming that the original label should be correctly annotated, we replaced 10% of the labels in the CoNLL2003 training dataset with different labels that do not conflict with the IOB2 format (the detail of the replacement is provided in the Appendix D). As a result, 3,329 sentences including pseudo-incorrect labels, referred to as replaced, and 5,356 sentences without replacements of annotations, referred to as unreplaced are generated by this method from 8,685 sentences.
We assign weights to this modified dataset using CrossWeigh and SubRegWeigh [^7] The hyperparameter settings are the same as those described in Section §4.3 and Appendix C.. Then we compared the average weights between the replaced and the original samples to confirm whether annotation error detection is correct. Here, we calculate the weights as (number of the same inferences as labels) / (number of inferences) instead of $w$ in Eq. (1). This is because we consider that the calculation using the minimum weight $w_{min}$ and the weight correction in CrossWeigh could have a negative effect on the fair analysis.
In Table 2, $\overline{w}_{incor}$ is the averaged weights assigned to the samples with pseudo-incorrect labels, where we expect them should be lower weights as incorrect labels. Similarly, $\overline{w}_{cor}$ is the averaged weights assigned to the original labels, where higher weights should be assigned as the correct labels.
@@ 5.2 Numbers of Tokenization Candidates
Table 3: Difference in the speed and the performance against three options of $K$.
In the proposed method, $K$ is an important hyperparameter affecting both speed and performance. We investigated the difference in the time for the annotation weighing and the $F_{1}$ score of the final model with three options of the number of tokenization candidates $K$ in the NER dataset. We examined the difference in the range of $K={10,50,500}$ with RoBERTa. The dropout rate for BPE-Dropout was $p=0.1$. Since $K$-means and Cos-Sim are techniques to reduce the variations in subword segmentation, a large value like $K=500$ makes them almost indistinguishable from the Random method. Therefore, only the Random method was investigated for $K=500$.
The results are shown in Table 3. Comparing the results between the ones with $K=10$ and $K=50$, one can see that $K$ does not have a large effect on the $F_{1}$ scores. However, the larger $K$ remarkably increases the time for the annotation weighing. For $K=500$, the $F_{1}$ score improved compared to Random with $K=10$ and $K=50$ on the CoNLL and CoNLL-CW datasets. This suggests that using a large number of subword sequences can effectively weigh annotation errors. However, such much large $K$ damages the speed of annotation weighing, making it a trade-off between the speed vs. the performance. While increasing $K$ to 500 in the Random method showed performance improvement, $K$-means exhibited the highest performance even with smaller $K$. This indicates that $K$-means can select sufficiently diverse subword sequences even with a small $K$.
@@
Table 4: Difference in the performance when using different models for scouting and final model.
@@ 5.3 Weighting with Different Backbone
In the experiment shown in §4, we used the same backbone for the scouting model $M$ for annotation weighing and the final model $M^{\prime}$. Herein, we investigated the performance difference when using different backbones in the NER dataset. Specifically, we examined the effect in $F_{1}$ scores when $M$ was RoBERTa and $M^{\prime}$ was LUKE. Similarly, we examined the case where $M$ was LUKE and $M^{\prime}$ was RoBERTa. Both models were trained with the hyperparameters explained in Table 7. We used $K$-means for selecting the tokenization candidates because it showed the best performance in the main results (§4).
@@ 6 Conclusion
We proposed SubRegWeigh, a method for annotation weighing using subword regularization, which offers faster annotation weighing compared to existing methods. In particular, subword sequence selection using $K$-means was four to five times faster than CrossWeigh for annotation weighing while contributing to better model performance than weighting with inference on all the large number of generated subword sequences.
@@
In this paper, we proposed a method for annotation weighing using subword regularization. However, since a deep learning model is used for error weighing, the calculated weights are not always guaranteed to be appropriate, and there is no assurance that all errors are weighed.
@@
Experiments presented in this work used datasets from previously published research Tjong Kim Sang and De Meulder (2003); Wang et al. (2019); Liu and Ritter (2023); Socher et al. (2013). These datasets were used to study NER or text classification models, which is consistent with their intended use. Because this study focuses on efficiently detecting annotation errors in the dataset, its potential risks and negative impact on society appear to be minimal. However, as indicated in the Limitation section, its output results may contain bias.
@@ Appendix A Example For Each Selection Method
Table 6: Examples of subword series by each selection method. Subword breaks are shown in the blanks.
In §3.2, we use 3 subword sequence selection methods: Random, Cos-Sim, and K-means. We show specific examples of the subword series obtained by each selection method in Table 6.
@@ Appendix C Detail About Model Setting
Table 7: model hyperparameter. *In $\mathrm{RoBERTa}$, the scouting model was trained in 5 epochs and the final model was trained in 20 epochs. Other hyper-parameters are the same in the scouting and final model.
@@
Appendix D Pseudo-Incorrect Label Method

### v2

@@ 4.1 Dataset
We briefly overview the dataset used in the experiments here. Appendix B also explains in detail.
@@
We used CoNLL-2003 Tjong Kim Sang and De Meulder (2003) with the official split for the training, validation, and test. The label format was changed from IOB1 to IOB2 Tjong Kim Sang and Veenstra (1999). In addition, we used CoNLL-CW Wang et al. (2019) and CoNLL-2020 Liu and Ritter (2023) for the test dataset. [^3] Because both of these dataset names are CoNLL++, we call CoNLL-CW and CoNLL-2020 to distinguish them.
@@
We used SST-2 Socher et al. (2013). The official training split was used for the training dataset and the validation split was used for the test dataset.
@@ 4.2 Compared Methods
For the baselines, we selected Vanilla and CrossWeigh. Vanilla is the final model that is trained on the original training split without any annotation weighing. CrossWeigh is the final model that is trained with the dataset weighed by the official code [^4] https://github.com/ZihanWangKi/CrossWeigh of the existing work Wang et al. (2019). Although CrossWeigh is specialized to NER, we forcibly utilize it for the text classification setting without Entity Disjoint Wang et al. (2019).
We evaluated the three options of our method introduced in §3.2, which are represented as SubRegWeigh (Random), SubRegWeigh (Cos-Sim), and SubRegWeigh ($K$-means), respectively. We used these proposed methods to weight the training dataset and evaluate the final model trained with this weighted dataset as well as CrossWeigh.
@@ 4.3 Model Settings
We employed $\mathrm{RoBERTa_{LARGE}}$ as the pre-trained language model for both NER and text classification. Besides $\mathrm{LUKE_{LARGE}}$ Yamada et al. (2020) is used for the NER setting because LUKE is an architecture for entity-related tasks. In this experimental setting, we used the same backbone model for both the scouting model and the final model555§5.3 discusses the case where using different pre-trained models between the scouting and final models.. For example, when RoBERTa is used in the scouting model, the final model will also use RoBERTa. The detailed model settings such as hyperparameters are provided in the Appendix C.
We set the three hyperparameters of CrossWeigh: the number of folds in mistake estimation $K=10$, the number of iterations of mistake estimation $T=3$, and the weight scaling factor $\epsilon=0.7$. BPE-Dropout Provilkov et al. (2020) was used to obtain the multiple tokenization candidates for SubRegWeigh because both RoBERTa and LUKE employ BPE for their tokenizers. BPE-Dropout has a hyperparameter $p$, where a higher $p$ results in finer tokenization (i.e., more different from the original tokenization). In the experiments, $p$ was set to $0.1$ unless otherwise noted. Additionally, by default, the hyperparameters for SubRegWeigh were set to $N=500$, $K=10$, and $w_{\mathrm{min}}=1/3$.
@@ 4.4 Evaluation Metrics
We evaluate the methods from the viewpoints of speed and performance. On the speed side, we measured the time for detecting the annotation errors and generating the weighted data for the training dataset in NER [^6] We measured the processing time with Xeon Platinum 8167M + NVIDIA V100 for RoBERTa and Xeon Gold 6230R + NVIDIA A100 for LUKE.. On the performance side, we measured the $F_{1}$ score in NER and the accuracy in text classification on each test dataset for the final models trained on the weighted training data.
@@
Table 1: The processing time and the performance of the final models with the weighted dataset. The best performing scores are highlighted in bold. and †† is SoTA score from Wang et al. (2021b) and  Zhou and Chen (2021).
@@ 4.5 Experimental Results
The results of our experiments are shown in Table 1. We reported the averaged scores and the standard deviations over the five independent runs.
4.5.1 Processing Time
The column “Processing time” shows the time taken to detect the annotation errors and assign weights for each sample in the training dataset in NER. As shown in this column, the proposed method, SubRegWeigh (Random), was maximum of over eight times faster than CrossWeigh. Even the most time-consuming method of SubRegWeigh ($K$-means) was five times faster with RoBERTa and four times faster with LUKE than CrossWeigh. The method with the random selection is the fastest among the proposed methods because it does not require any of the calculations about TF-IDF, cosine similarity, and $K$-means clustering. In addition, in the random method, only $K(=N=10)$ tokenization candidates were generated, whereas the other methods of SubRegWeigh need to select $K$ candidates from the $N(=500)$-sampled tokenization candidates, which results in longer processing time. Although the proposed method of Cos-Sim and $K$-means takes longer time than Random, it is remarkably faster than the existing method for the automatic annotation weighing, CrossWeigh.
@@ 4.5.2 Performance on NER
The performance on CoNLL-CW is the most important in our evaluation because annotation errors in the test split are removed, which indicates that we can compare the methods on the most unpolluted dataset [^7] The vanilla baseline score of LUKE exceeds the SoTA score Zhou and Chen (2021) for CoNLL-CW, which we consider because of the successful hyperparameter setting.. One can see that SubRegWeigh ($K$-means) achieves higher performance than the vanilla method on the test dataset of CoNLL-CW. Specifically, SubRegWeigh ($K$-means) improved the $F_{1}$ scores by 0.18 points with RoBERTa and 0.20 points with LUKE compared to each vanilla baseline. This result indicates that the proposed method of annotation weighing contributes to the performance improvement in the clean test dataset. Furthermore, SubRegWeigh ($K$-means) achieves higher scores than CrossWeigh, which demonstrates that the proposed method is a reasonable alternative to the existing method.
Both SubRegWeigh and CrossWeigh scored lower than Vanilla in the CoNLL valid and test datasets, which contain annotation errors. These results indicate the remarkable negative effect of the annotation errors in the valid and test datasets and the difficulty of evaluating the models appropriately.
Among the proposed methods, Random scores lower than $K$-means in many cases. We consider this because the random selection of tokenization candidates leads to the selection of similar subword sequences, causing a bias in the inference results. Cos-Sim also scored lower than $K$-means in all cases, which shows that the naive method of selecting tokenization candidates does not contribute to performance improvement. We consider that the $K$-means clustering can handle the diverse range of subword tokenization candidates, and this leads to the successful result.
@@ 4.5.3 Performance on Text Classification
For text classification, SubRegWeigh ($K$-means) achieved the highest accuracy, and SubRegWeigh (random) had lower accuracy, similar to the results in NER. One reason for the lack of performance improvement with CrossWeigh in SST-2 is that Entity Disjoint is important for CrossWeigh but unavailable in the text classification task.
From the results on the NER and text classification datasets, we conclude that the proposed method is superior to the existing method in terms of processing time and task performance. Especially, we can say SubRegWeigh ($K$-means) is the best option, which achieves the highest task performance with approximately four to five times faster than CrossWeigh.
@@ 5.1 Pseudo-incorrect Label Test
$\overline{w}_{\mathrm{cor}}$ $\overline{w}_{\mathrm{incor}}$
@@
Table 2: Averaged weights and performance ($F_{1}$) for the pseudo-incorrect dataset. $\overline{w}_{\mathrm{cor}}$ and $\overline{w}_{\mathrm{incor}}$ are the averaged weights for the samples with original labels and ones with the pseudo-incorrect labels, respectively.
We verified whether SubRegWeigh can accurately detect annotation errors in the dataset. We evaluate this capability using a modified dataset with some labels flipped to incorrect labels artificially.
Assuming that the original label should be correctly annotated, we replaced 10% of the labels in the CoNLL2003 training dataset with different labels that do not conflict with the IOB2 format (Appendix D). Then, 3,329 sentences including pseudo-incorrect labels, and 5,356 sentences with original annotations are obtained.
We assign weights to this modified dataset using CrossWeigh and SubRegWeigh [^8] The hyperparameter settings are the same as those described in Section §4.3 and Appendix C.. Then we compared the averaged weights between the pseudo-incorrect and the original samples to confirm whether annotation error detection is correct. Here, we calculate the weights as $\frac{C}{K}$, where $C$ is the number of inferences that the model predicts the original label, instead of $w$ in Eq. (1). This is because we consider that the minimum weight $w_{min}$ and the weight correction in CrossWeigh could have a negative effect on the fair analysis.
In Table 2, $\overline{w}_{\mathrm{incor}}$ is the averaged weights assigned to the samples with pseudo-incorrect labels, where we expect them should be lower as incorrect labels. Similarly, $\overline{w}_{\mathrm{cor}}$ is the averaged weights assigned to the original labels, where higher weights should be assigned as the correct labels.
@@ 5.2 Numbers of Tokenization Candidates
Table 3: Difference in the processing time (hh:mm) and the performance ($F_{1}$) against three options of $K$.
@@
Table 4: Difference in the performance ($F_{1}$) when using different pre-trained models for scouting and final model.
@@
$K$ is an important hyperparameter affecting both processing time and performance of SubRegWeigh. We investigated the difference in the processing time and the $F_{1}$ score of the final model with three options of the number of tokenization candidates $K$ in the NER dataset. We examined the difference in the range of $K={10,50,500}$ with RoBERTa. The dropout rate for BPE-Dropout was $p=0.1$. Since $K$-means and Cos-Sim are techniques to reduce the variations in subword segmentation, a large value like $K=500$ makes them almost indistinguishable from the Random method. Therefore, only the Random method was investigated for $K=500$.
The results are shown in Table 3. Comparing the results between the ones with $K=10$ and $K=50$, one can see that $K$ does not have a large effect on the $F_{1}$ scores. However, the larger $K$ remarkably increases the processing time. For $K=500$, the $F_{1}$ score improved compared to Random with $K=10$ and $K=50$ on the CoNLL and CoNLL-CW datasets. This suggests that using a large number of subword sequences can effectively weigh annotation errors. However, such much large $K$ damages the speed of annotation weighing, making it a trade-off between the speed vs. the performance. While increasing $K$ to 500 in the Random method showed performance improvement, $K$-means exhibited the highest performance even with smaller $K$. This indicates that $K$-means can select sufficiently diverse subword sequences.
@@ 5.3 Weighting with Different Backbone
In the experiment shown in §4, we used the same backbone for the scouting model $M$ for annotation weighing and the final model $M^{\prime}$. Herein, we investigated the performance difference when using different backbones in NER. Specifically, we examined the effect in $F_{1}$ scores when $M$ was RoBERTa and $M^{\prime}$ was LUKE. Similarly, we examined the case where $M$ was LUKE and $M^{\prime}$ was RoBERTa. Both models were trained with the hyperparameters explained in Table 7. We used $K$-means for selecting the tokenization candidates because it showed the best performance in the main results (§4).
@@ 6 Conclusion
We proposed SubRegWeigh, a method for annotation weighing using subword regularization, which offers faster annotation weighing than existing methods. In particular, subword sequence selection using $K$-means was four to five times faster than CrossWeigh for annotation weighing and contributed to better performance than weighting with all of the large number of subword sequences.
@@
In this paper, we proposed a method for annotation weighing using subword regularization. However, since a deep learning model is used for error weighing, the calculated weights are not always guaranteed to be appropriate.
As mentioned in 5.4, if there is not enough variation of the subword sequence, such as uppercase-only sentences, our proposed method often assigns incorrect weights.
In §5.1, we created pseudo-incorrect labels by randomly replacing the original labels with incorrect labels. This error distribution may differ from real-world annotation error distribution.
@@
Experiments presented in this work were performed with existing datasets Tjong Kim Sang and De Meulder (2003); Wang et al. (2019); Liu and Ritter (2023); Socher et al. (2013). These datasets were used to study NER or text classification models, which is consistent with their intended use. Because this study focuses on efficient annotation error detection in the dataset, its potential risks and negative impact on society appear to be minimal.
@@ Appendix A Example For Each Selection Method
Table 6: Examples of subword sequences by each selection method. Subword breaks are shown in the blanks.
In §3.2, we use 3 subword sequence selection methods: Random, Cos-Sim, and K-means. We show specific examples of the subword sequences obtained by each selection method in Table 6.
@@ B.1 NER
Figure 2: The overview of SubRegWeigh in the case of NER.
The overview in the case of NER is shown in Figure 2.
@@ Appendix C Detail About Model Setting
Table 7: Model hyperparameters. *In $\mathrm{RoBERTa}$, the scouting model was trained in 5 epochs and the final model was trained in 20 epochs. Other hyper-parameters are the same in the scouting and final model.
@@
Appendix D Pseudo-Incorrect Labeling Method